OpenStudy (anonymous):

What is the quadratic equation with the roots -3+2i?

Nnesha (nnesha):

imaginary solutions if one of them is -3+2i then 2nd one would be -3-2i remember when we take square root we get plus/minus solution so x =-3+2i x=-3-2i set it equal to zero \[\huge\rm (x-(-3+2i))(x-(-3-2i))\]

Nnesha (nnesha):

now first distribute by negative